OpenStudy (anonymous):

solve the logarithmic equation log7 x +log 7 (x-48)=2

Nnesha (nnesha):

\[\huge\rm log_7 x + \log_7 (x -48) =2 \] what rule of log you should apply first ?? if there is positive sign how would you condense that ??

OpenStudy (anonymous):

I know first step combine logs because they are the same and log7(x(x^2-48))=2

Nnesha (nnesha):

gO_Od!!!

OpenStudy (anonymous):

log7(x^2-48X)=2

Nnesha (nnesha):

when you multiply variables you should add their exponents \[\huge\rm x^m \times x^n = x^{m+n}\]

OpenStudy (anonymous):

here is where I need help I know you have to bring over the 2 but I forget what about log 7

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @sam87211 log7(x^2-48X)=2 \(\color{blue}{\text{End of Quote}}\) this one distribute both term by x so......

Nnesha (nnesha):

nvm i thought there is (x^2 -48)

Nnesha (nnesha):

now you have to change log form to exponential